# On the Parallelizability of Approval-Based Committee Rules

### v1

@@ 5 Conclusion and Future Work
We systematically studied the parallelizability of finding a winning committee for all of the polynomial-time Approval-Based Committee rules studied by Lackner and Skowron (2023) in their recent textbook. We find that with the exception of two simple cases, as pointed out by Lackner and Skowron (2023), all of the remaining rules are inherently sequential. We further explored the parallelizability of ABC rules by considering restricted domains and found that for the natural settings of single-peaked and single-crossing votes finding a Chamberlin-Courant committee is parallelizable. We showed this result by giving algorithms that show these problems in the complexity class OptL, the first results of this type in the computational social choice.
@@ Acknowledgments
This work was supported in part by NSF grants CCF-2421977 and CCF-2421978.
@@ Appendix A Logspace Algorithms for AV and SAV
For the sake of completeness—to ensure we study all polynomial-time rules in (Lackner and Skowron, 2023), we show in this section that Approval Voting and Satisfaction Approval Voting are in FL (the class of functions computable in logarithmic space).
@@
2:     count votes for $c$ and store in $cvotes$
@@
5:        count votes for $d$ and store in $dvotes$
6:        if  $cvotes>dvotes$ or ($cvotes=dvotes$ and $c<d$then
7:           $rank$ += $1$
@@ C.1 Sequential Thiele Methods
Let $w:\mathbb{Z}_{\geq 0}\xrightarrow{}\mathbb{R}$ be a fixed non-decreasing function. Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. seq-$w$-Thiele sequentially builds a committee, $W$, of size $k$ by repeatedly adding to $W$ the candidate whose inclusion increases the value
@@
[…]
@@
Computing the winning committee for seq-$w$-Thiele is inherently sequential for any nondecreasing function $w$ with $w(0)=0$, $w(1)=1$, and $w(2)=1+\epsilon$ for any real $\epsilon\in[0,1)$.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by seq-$w$-Thiele includes $c$ if and only if $v$ is included in the LFMIS of $G$.
We add to $C$ each vertex in $G$ as well as $m$ extra candidates, labeled $m+1$ to $2m$. We add a vote to $V$ for each edge in $G$ as well as three singleton votes for each extra candidate. Let $k=m$ and $c=v$.
We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. Note that every candidate in $C$ has three votes. Thus, adding any vertex from $C$ will result in increasing our score to 3. Moreover, each voter in $V$ votes for at most two candidates, so only the values $w(0),w(1),$ and $w(2)$ are relevant to our reduction.
The first candidate picked corresponds to vertex 1 due to lexicographic tie-breaking. Let $u$ be the candidate picked in the next round. Observe that if $u$ and $1$ share a vote in $V$ (i.e., $u$ is adjacent to $1$ in $G$), then adding $u$ to the committee will increase the score by $31+1+\epsilon<$ 3. Whereas, any $u$ that does not share a vote with $1$ will increase the score by 3. Thus, $u$ must not be adjacent to $1$. Generalizing this argument, we obtain an increase in score of $2+\epsilon$, $1+2\epsilon$, or $3\epsilon$ if a candidate sharing a vote with one, two, or three other candidates is added. Thus, the seq-$w$-Thiele procedure will always prefer adding a candidate that does not share a vote with any previously added candidates (i.e., a vertex in $G$ that is not adjacent to any previously added vertex).
From this observation and lexicographic tie-breaking, it is easy to see that the candidates corresponding to the LFMIS of $G$ will be added to the committee first. Once no more candidates from $V(G)$ can be added, we will start picking candidates from $m+1$ to $2m$ until the algorithm terminates.
@@
Let $w:\mathbb{Z}_{\geq 0}\xrightarrow{}\mathbb{R}$ be a fixed non-decreasing function. Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. rev-seq-$w$-Thiele builds a committee, $W$, of size $k$ by first adding every candidate to $W$ and then repeatedly removing from $W$ the candidate whose removal decreases the value
@@
Computing the winning committee for rev-seq-$w$-Thiele is inherently sequential for any nondecreasing function $w$ with $w(0)=0$, $w(1)=1$, and $w(2)=1+\epsilon$ for any real $\epsilon\in[0,1)$.
@@
Suppose we are given an instance $(G,v)$ of $\overline{\text{LFMIS}}$, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by rev-seq-$w$-Thiele does not include $c$ if and only if $v$ is included in the LFMIS of $G$.
We add to $C$ each vertex in $G$ as well as $2m$ extra candidates, labeled $m+1$ to $3m$. We add a vote to $V$ for each edge in $G$ as well as three votes for every two extra candidates, i.e., three votes $\{m+1,m+2\}$, three votes $\{m+3,m+4\}$, $\ldots$, and three votes $\{3m-1,3m\}$. Finally, let $k=m$ and $c=v$.
We now show that the candidates removed from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. Note that every candidate in $C$ belongs to exactly three votes. Since each voter votes for two candidates, the only values of $w$ relevant to us at $w(0)$, $w(1)$, and $w(2)$. Recall that initially all candidates are included in the winning committee $W$. Observe that removing any candidate from $W$ will decrease the score by $3\epsilon$. The first candidate removed corresponds to vertex 1 due to lexicographic tie-breaking. Let $u$ be the candidate picked in the next round. Observe that if $u$ and $1$ share a voter in $V$ (i.e., $u$ is adjacent to $1$ in $G$), then removing $u$ from the committee will decrease the score by $31+2\epsilon<$ 3. Whereas, any $u$ that does not share a vote with $1$ will decrease the score by $3\epsilon$. Thus, $u$ must not be adjacent to $1$. Generalizing this argument, we obtain a decrease in score of $3\epsilon$, $1+2\epsilon$, or $2+\epsilon$ if a candidate sharing a vote with one, two, or three other previously removed candidates is removed.
Thus, the rev-seq-$w$-Thiele procedure will always prefer removing a candidate that does not share a vote with any previously removed candidates (i.e., a vertex in $G$ that is not adjacent to any vertex previously added to the LFMIS).
From this observation and lexicographic tie-breaking, it is easy to see that the candidates corresponding to the LFMIS of $G$ will be removed from the committee first. Once no more candidates from $V(G)$ can be removed without decreasing the score more than $3\epsilon$, we will start picking the extra candidates labeled $m+1$ to $3m$ until the algorithm terminates; the candidates removed from the extra candidates will also decrease the score by $3\epsilon$.
Note that $k$ is large enough (at least the number of vertices in $G$) so that every candidate corresponding to a vertex in the LFMIS of $G$ will be removed. Similarly, the number of extra candidates is large enough so that we are not forced to remove any candidates corresponding to the vertices in $G$ after the LFMIS candidates have been picked.
@@
Computing the winning committee for seq-Phragmén is inherently sequential.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by seq-Phragmén includes $c$ if and only if $v$ is included in the LFMIS of $G$.
We add to $C$ each vertex in $G$ as well as $m$ extra candidates, labeled $m+1$ to $2m$. We add a vote to $V$ for each edge in $G$ as well as three singleton votes for each extra candidate. Let $k=m$ and $c=v$.
We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. Recall that $G$ is 3-regular and the candidates labeled $m+1$ to $2m$ have exactly three votes, so, $\|N(c)\|=3$ for all $c\in C$. Thus, we are only concerned with minimizing the load
$\ell_{r}(c)=1+\sum_{v\in N(c)}y_{r-1}(v)$
in each round. The first candidate picked corresponds to vertex 1 due to lexicographic tie-breaking.
We have $\ell_{1}(1)=1/3$, and thereby $y_{r}(i)=1/3$ for each $i\in N(1)$. Let $u$ be the candidate picked in the next round. Observe that if $u$ and $1$ share a vote in $V$ (i.e., $u$ is adjacent to $1$ in $G$), then we obtain a maximum load of $\ell_{2}(u)=(1+1/3)/3=4/9$. Whereas, any $u$ that does not share a vote with $1$ will increase the maximum load by $\ell_{2}(u)=(1+0)/3=1/3$. Thus, $u$ must not be adjacent to $1$.
Generalizing this argument, we obtain an increase in load of $1/3$ whenever a candidate that does not share a vote with any previously selected candidate, but a larger load otherwise. Thus, the seq-Phragmén procedure will always prefer adding a candidate that does not share a vote with any previously added candidates (i.e., a vertex in $G$ that is not adjacent to any previously added vertex).
From this observation and lexicographic tie-breaking, it is easy to see that the candidates corresponding to the LFMIS of $G$ will be added to the committee first. Once no more candidates from $V(G)$ can be added, we will start picking the candidates labeled $m+1$ to $2m$ until the algorithm terminates.
@@ C.3 Greedy Monroe
For the case where $n$ is not divisible by $k$, the upper bound used for the number of selected voters each round is set like so. Let $d=n\mod k$. For rounds $1$ to $d$, $\left\lceil n/k\right\rceil$ voters are selected. For rounds $d$ to $k$ $\left\lfloor n/k\right\rfloor$ voters are selected.
@@
Computing the winning committee for Greedy Monroe is inherently sequential.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by GM includes $c$ if and only if $v$ is included in the LFMIS of $G$.
We add to $C$ each vertex in $G$ as well as $m$ extra candidates, labeled $m+1$ to $2m$. We add a vote to $V$ for each edge in $G$ as well as three singleton votes for each extra candidate. Let $k=m$ and $c=v$.
We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. Observe that
$\frac{n}{k}=\frac{3m/2+3m}{m}=3/2+3\geq 3$
Recall that $G$ is 3-regular and the candidates labeled $m+1$ to $2m$ have exactly three votes. So, whenever a candidate is selected, it is selected as the representative as all of its voters.
Since each candidate has the same number of votes, the first candidate picked corresponds to vertex 1 due to lexicographic tie-breaking.
Candidate 1 is chosen as the representative of all of its voters, i.e., every voter approving of candidate 1 is added to $G_{1}$. Thus, $V_{2}$, the unrepresented voters of the next round, now include all voters except those that approved candidate 1.
Observe that any candidate that shared a vote with candidate 1 now has two voters approving them, while any candidate not sharing a vote with candidate 1 has three voters approving them. Generalizing this argument, every round we prefer selecting a candidate that does not share a vote with any previously selected candidate (i.e., a vertex in $G$ that is not adjacent to any previously added vertex).
From this observation and lexicographic tie-breaking, it is easy to see that the candidates corresponding to the LFMIS of $G$ will be added to the committee first. Once no more candidates from $V(G)$ can be added, we will start picking the candidates labeled $m+1$ to $2m$ until the algorithm terminates.
@@
C.4 Method of Equal Shares

### v3

@@ 5 Conclusion and Future Work
We systematically studied the parallelizability of finding a winning committee for all of the polynomial-time Approval-Based Committee rules studied by Lackner and Skowron (2023) in their recent textbook. We find that with the exception of two simple cases, as pointed out by Lackner and Skowron (2023), all of the remaining rules are inherently sequential. We further explored the parallelizability of ABC rules by considering restricted domains and found that for the natural settings of single-peaked and single-crossing votes finding a Chamberlin-Courant committee is parallelizable. We showed this result by giving algorithms that show these problems are in the complexity class OptL, the first results of this type in computational social choice.
@@ Acknowledgments
We would like to thank Dominik Peters, Tomoyuki Yamakami, and the anonymous reviewers for their helpful comments and suggestions. This work was supported in part by NSF grants CCF-2421977 and CCF-2421978.
@@ Appendix A Logspace Algorithms for AV and SAV
For the sake of completeness—to ensure we study all polynomial-time rules in Lackner and Skowron (2023), we show in this section that Approval Voting and Satisfaction Approval Voting are in FL (the class of functions computable in logarithmic space).
@@
2:  count votes for $c$ and store in cvotescvotes
@@
5:   count votes for $d$ and store in dvotesdvotes
6:   if  $cvotes>dvotes$ or ($cvotes=dvotes$ and c<dc<dthen
7:    rankrank += 11
@@ C.1 Sequential Thiele Methods
Let $w:\mathbb{Z}_{\geq 0}\xrightarrow{}\mathbb{R}$ be a fixed nondecreasing function with $w(0)=0$. Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. seq-$w$-Thiele sequentially builds a committee, $W$, of size $k$ by repeatedly adding to $W$ the candidate whose inclusion increases the value
@@
Another popular Thiele method is Proportional Approval Voting (PAV). In PAV, a voter’s contribution to the score of the committee is based on how many approved candidates it voted for; if it voted for one candidate that was approved, it contributes a score of $1$, if it voted for 2, it contributes a score of $1+1/2$, and so on. This scoring is captured by the harmonic function: $h(x)=\sum_{k=1}^{x}1/x$. In Lackner and Skowron (2023), seq-$h$-Thiele is referred to as seq-PAV. Below, we prove that computing seq-$w$-Thiele is inherently sequential for any rule where we have $w(0)=0$, $w(1)=1$, and $w(2)=1+\epsilon$, where $\epsilon\in[0,1)$ is a rational number. Note that seq-CC and seq-PAV both fit this class of Thiele methods.
@@
Computing the winning committee for seq-$w$-Thiele is inherently sequential for any nondecreasing function $w$ with $w(0)=0$, $w(1)=1$, and $w(2)=1+\epsilon$ for any rational $\epsilon\in[0,1)$, even when tie-breaking is never invoked.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected includes $c$ if and only if $v$ is included in the LFMIS of $G$.
Let $C$ have a candidate for each vertex of $G$. Let $p:=\lceil\frac{4}{1-\epsilon}\rceil$. For each $\{a,b\}\in E(G)$, we add $pm$ votes $\{a,b\}$ to $V$. We also add (ma)(m-a) singleton votes for each $a\in C$. Note that each candidate $a\in\{1,2,\ldots,m\}$ has 3pm+(ma)3pm+(m-a) votes. Finally, add $m$ extra candidates, labeled $m+1$ to $2m$, that have $3pm-1,3pm-2,\ldots,3pm-m$ votes. Note that these extra candidates have fewer votes than all the candidates corresponding to vertices in $G$. It is clear that there are no ties in the election. Finally, we set $c=v$ and $k=m$. We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$.
Claim 24.
Let $a$ and $b$ be unelected candidates in round $r+1$. If $b$ shares votes with a previously elected candidate and $a$ does not, then adding $a$ to $W$ increases the score more than adding $b$.
Proof.
Any candidate $b$ that shares votes with a previously elected candidate gives an increase in score of
$2pm+(m-b)+\epsilon pm\leq(2p+1+\epsilon p)m$
Any candidate $a$ that shares no votes with a previously elected candidate has at least 3pmm3pm-m votes.
We will show that 3pmm3pm-m is greater than $(2p+1+\epsilon p)m$. To do so, it is sufficient to show that
$3pm-m$ $>(2p+1+\epsilon p)m$
$pm$ $>2m+\epsilon pm$
$(p-\epsilon p)m$ $>2m$
$p-\epsilon p$ $>2$
Since $x+1>\lceil x\rceil\geq x$ for any $x$, we have:
$p-\epsilon p=\left\lceil\frac{4}{1-\epsilon}\right\rceil-\epsilon\left\lceil\frac{4}{1-\epsilon}\right\rceil\geq\frac{4}{1-\epsilon}-\epsilon\left(\frac{4}{1-\epsilon}+1\right)=4-\epsilon$
We have $4-\epsilon>2\implies 2>\epsilon$ by definition of $\epsilon$. ∎
By the claim above and the fact that candidates with more votes (i.e., vertices with earlier lexicographic order) are preferred, it is clear that the candidates picked will correspond exactly to the vertices in the LFMIS of $G$. It is also clear that the tie-breaking rule is never invoked while building the committee.
Once adding candidates from $V(G)$ is suboptimal, we will start picking the candidates labeled $m+1$ to $2m$ until the algorithm terminates.
@@
Let $w:\mathbb{Z}_{\geq 0}\xrightarrow{}\mathbb{R}$ be a fixed nondecreasing function with $w(0)=0$. Suppose we are given candidates $C$, collection of votes $V$, and committee size $k$. rev-seq-$w$-Thiele builds a committee, $W$, of size $k$ by first adding every candidate to $W$ and then repeatedly removing from $W$ the candidate whose removal decreases the value
@@
Computing the winning committee for rev-seq-$w$-Thiele is inherently sequential for any nondecreasing function $w$ with $w(0)=0$, $w(1)=1$, and $w(2)=1+\epsilon$ for any rational $\epsilon\in[0,1)$, even when tie-breaking is never invoked.
@@
Suppose we are given an instance $(G,v)$ of $\overline{\text{LFMIS}}$, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by rev-seq-$w$-Thiele does not include $c$ if and only if $v$ is included in the LFMIS of $G$.
Let $C$ have a candidate for each vertex of $G$. Let $p:=\lceil\frac{4}{1-\epsilon}\rceil$. For each $\{a,b\}\in E(G)$, we add $pm$ votes $\{a,b\}$ to $V$. We also add $a$ singleton votes for each $a\in C$. Note that each candidate $a\in\{1,2,\ldots,m\}$ has $3pm+a$ votes. Finally, add $2m$ extra candidates, labeled $m+1$ to $3m$.
We add $3pm$ votes $\{m+1,m+2\}$, $3pm$ votes $\{m+3,m+4\}$, $\ldots$, and $3pm$ votes $\{3m-1,3m\}$. We also add $a$ singleton votes for each $a\in\{m+1,m+2,\ldots,3m\}$. So each extra candidate $a$ has $3pm+a$ votes. Finally, let k=3mm=2mk=3m-m=2m and $c=v$.
We now show that the candidates removed from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$.
Claim 27.
Let $a$ and $b$ be selected candidates in round $r+1$. If $b$ shares votes with a previously removed candidate and $a$ does not, then removing $a$ from $W$ decreases the score less than removing $b$.
Proof.
Removing $b$ gives a decrease of
$\epsilon(2pm)+pm+b\geq\epsilon(2pm)+pm$
Removing $a$ gives a decrease of
$\epsilon(3pm)+a\leq\epsilon(3pm)+3m$
We will show that $\epsilon(3pm)+3m<\epsilon(2pm)+pm$. To do so, it is sufficient to show that
$\epsilon(2pm)+pm$ $>\epsilon(3pm)+3m$
$pm-\epsilon pm$ $>3m$
$p-\epsilon p$ $>3$
Since $x+1>\lceil x\rceil\geq x$ for any $x$, we have:
$p-\epsilon p=\left\lceil\frac{4}{1-\epsilon}\right\rceil-\epsilon\left\lceil\frac{4}{1-\epsilon}\right\rceil\geq\frac{4}{1-\epsilon}-\epsilon\left(\frac{4}{1-\epsilon}+1\right)=4-\epsilon$
We have $4-\epsilon>3\implies 1>\epsilon$ by definition of $\epsilon$. ∎
By the claim above and the fact that candidates with fewer votes (i.e., vertices with earlier lexicographic order) are preferred, it is clear that the candidates removed will correspond exactly to the vertices in the LFMIS of $G$. It is also clear that the tie-breaking rule is never invoked while building the committee.
Once removing candidates from $V(G)$ is suboptimal, we will start removing the candidates labeled $m+1$ to $3m$ until the algorithm terminates.
Note that $k$ is set so that $m$ candidates will be removed, i.e., every candidate corresponding to a vertex in the LFMIS of $G$ will be removed. Similarly, the number of extra candidates is large enough so that we are not forced to remove any candidates corresponding to the vertices in $G$ after the LFMIS candidates have been picked.
@@
Computing the winning committee for seq-Phragmén is inherently sequential, even when tie-breaking is never invoked.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. Without loss of generality, we will assume that $m\geq 3$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by seq-Phragmén includes $c$ if and only if $v$ is included in the LFMIS of $G$.
Our construction is similar to the proof for Theorem 12. Let $C$ have a candidate for each vertex of $G$. For each $\{a,b\}\in E(G)$, we add $m^{2}$ votes $\{a,b\}$ to $V$. We also add (ma)(m-a) votes for each $a\in C$. Note that each candidate $a\in\{1,2,\ldots,m\}$ has $\|N(a)\|=3m^{2}+(m-a)$ votes. Finally, add $m$ extra candidates, labeled $m+1$ to $2m$, that have $3m^{2}-1,3m^{2}-2,\ldots,3m^{2}-m$ votes. It is clear that there are no ties in the election. Finally, we set $c=v$ and $k=m$.
We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. We first prove the following claim.
Claim 30.
Let $a$ and $b$ be unelected candidates in round $r$. If $b$ shares votes with an elected candidate and $a$ does not, then $\ell_{r}(b)>\ell_{r}(a)$.
Proof.
Note that we have
$\ell_{r}(a)=\frac{1+\sum_{i\in N(a)}y_{r-1}(i)}{\|N(a)\|}=\frac{1}{\|N(a)\|}\leq\frac{1}{3m^{2}-m}$
and Suppose $b$ shares votes with a previously elected candidate, $c$. The load of all the votes that $b$ and $c$ share is at least $1/\|N(c)\|$. There are $m^{2}$ such votes. Thus, we have:
$\ell_{r}(b)$ $=\frac{1+\sum_{i\in N(b)}y_{r-1}(i)}{\|N(b)\|}$
$\geq\frac{1+m^{2}\left(\frac{1}{\|N(c)\|}\right)}{\|N(b)\|}$
$\geq\frac{1+m^{2}\left(\frac{1}{3m^{2}+m}\right)}{3m^{2}+m}$
It can be verified that
$\frac{1+m^{2}\left(\frac{1}{3m^{2}+m}\right)}{3m^{2}+m}>\frac{1}{3m^{2}-m}$
for all $m\geq 3$. ∎
In the first round, the candidate with the most votes is preferred since it minimizes the maximum load. Thus, candidate $1$ is picked. By Claim 30 and the fact that candidates with more votes (i.e., vertices with earlier lexicographic order) are preferred, it is clear that the candidates picked after candidate 1 will correspond exactly to the vertices in the LFMIS of $G$.
Once adding candidates from $V(G)$ is suboptimal, we will start picking the candidates labeled $m+1$ to $2m$ until the algorithm terminates.
@@ C.3 Greedy Monroe
For the case where $n$ is not divisible by $k$, the upper bound used for the number of selected voters each round is set like so. Let $d=n\mod k$. For rounds $1$ to $d$, $\left\lceil n/k\right\rceil$ voters are selected. For rounds $d+1$ to $k$ $\left\lfloor n/k\right\rfloor$ voters are selected.
@@
Computing the winning committee for Greedy Monroe is inherently sequential, even when tie-breaking is never invoked.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by GM includes $c$ if and only if $v$ is included in the LFMIS of $G$.
Let $C$ have a candidate for each vertex of $G$. For each $\{a,b\}\in E(G)$, we add $2m$ votes $\{a,b\}$ to $V$. We also add (ma)(m-a) singleton votes for each $a\in C$. Note that each candidate $a\in\{1,2,\ldots,m\}$ has 6m+(ma)6m+(m-a) votes. Finally, add $m$ extra candidates, labeled $m+1$ to $2m$, that have $5m+1,5m+2,\ldots,6m$ votes. It is clear that there are no ties in the election. Finally, we set $c=v$ and $k=m$. We now show that the candidates picked from $V(G)$ correspond exactly to the vertices in the LFMIS of $G$. Observe that
$\frac{n}{k}=\frac{(\frac{3m}{2})2m+\sum_{a=1}^{m}(m-a)+\sum_{a=1}^{m}(5m+a)}{m}=9m$
Note that the maximum number of voters a candidate has is $6m+m-1\leq 9m$. Thus, whenever a candidate is elected, it is selected as the representative of all of its voters. We now prove the following claim.
Claim 33.
Let $a$ and $b$ be unelected candidates in round $r+1$. If $b$ shares votes with a previously elected candidate and $a$ does not, then $a$ has more votes in $V_{r+1}$ than $b$.
Proof.
Since a candidate is the representative of all of its voters and $b$ shares $2m$ candidates with a previously elected voter, $b$ has at most 6m+m12m=5m16m+m-1-2m=5m-1 votes in $V_{r+1}$. Any candidate $a$ that shares no votes with a previously elected candidate has at least $5m+1$ votes. ∎
By the claim above and the fact that candidates with more votes (i.e., vertices with earlier lexicographic order) are preferred, it is clear that the candidates picked will correspond exactly to the vertices in the LFMIS of $G$.
Once adding candidates from $V(G)$ is suboptimal, we will start picking the candidates labeled $m+1$ to $2m$ until the algorithm terminates.
@@
C.4 Method of Equal Shares + Phragmén